Circular No. 7105

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999X IN CGCG 180-22 M. Schwartz, Cottage Grove, OR, reports his discovery of an apparent

supernova (mag 16.1) on an unfiltered CCD frame taken on Jan. 23.23 UT utilizing the Tenagra I 0.35-m automated supernova patrol telescope in Sonoita, AZ, internet-controlled from Cottage Grove. SN 1999X is located at α = 8^{h}54^{m}32^{s}.23, δ = +36^{o}30′41′′.1 (equinox 2000.0), which is ∼ 4′′.1 east and 5′′.5 north of the nucleus of CGCG 180-22. The new object was confirmed on a frame taken on Jan. 27.13 at mag 15.3, showing significant brightening. Schwartz adds that the object does not appear on either the first or the second Palomar Sky Survey scans, and that it is absent from the USNO A1.0 catalogue.

P. Garnavich, S. Jha, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 1999X was obtained and reduced by P. Berlind (Fred L. Whipple Observatory) on Feb. 6.3 UT using the Tillinghast telescope. The spectrum shows a strong Si II absorption feature with a minimum at 629.6 nm, indicating that this is a type-Ia event. Other features suggest an age between 7 and 10 days past maximum light. The recession velocity of CGCG 180-22 is 7500 km/s (Ramella et al. 1995, A.J. 109, 1458), so that the supernova photospheric expansion velocity derived from the Si II line is 10 050 km/s.

RX J0052.1—7319 A. Udalski, Warsaw University Observatory, on behalf of the OGLE

microlensing team, communicates: “The proposed optical counterpart of RX J0052.1–7319 (cf. IAUC 7101) was identified in the OGLE photometric database (object SMC SC6 99923) as a long-term variable star with quasiperiodic light variation of amplitude 0.13 mag in the I band; the possible period of ∼ 600–700 days is uncertain because of comparable length with the current 745-day span of the OGLE SMC database (1997 Jan. 17-1999 Feb. 1). The mean photometry of the star is V = 14.667, B − V = −0.005, V − I = +0.146 (cf. OGLE photometric maps of the SMC, Udalski et al. 1998, Acta Astron. 48, 147). Photometry of the star is available from

the OGLE Internet archive at http://www.astrouw.edu.pl/~ftp/ogle.”

1999 February 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green